### v1

@@ D.2 Separating Simultaneous Treatment Effects
Under Assumption D.1, we can estimate the counterfactual series $Y_{1,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of Graves before the coming-out event as in Section 3.2. Thus, as shown in (D.2), the challenge in disentangling our causal effects of interest is to estimate $Y_{1,t}^{0,1}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved if Graves were already part of the LGB community prior to the 2022 LGBT Pride Month.
@@
Under Assumption (D.1), we observe $Y_{j,t}=Y_{j,t}^{0,1}$ for all $t>T^{pre}$, and we can estimate the counterfactual series $Y_{j,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of character $j$ before the beginning of the 2022 LGBT Pride Month. We can then estimate $\gamma_{t}^{L}$ by computing the differences between character $j$’s observed pick rates and the synthetic counterfactual for all $t>T^{pre}$:
@@
Under Assumption D.2, the relationship $\tau_{t}^{C}=\tau_{t}-\gamma_{t}^{L}$ holds. Thus, if the estimated effects of LGBT Pride Month on players’ preferences for LGB characters are small relative to the estimated effects of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4.1 must be primarily attributed to Graves’ disclosure.

### v3

@@ D.2 Separating Simultaneous Treatment Effects
Under Assumption 1, we can estimate the counterfactual series $Y_{1,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of Graves before the coming-out event as in Section 3.2. Thus, as shown in (D.2), the challenge in disentangling our causal effects of interest is to estimate $Y_{1,t}^{0,1}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved if Graves were already part of the LGB community prior to the 2022 LGBT Pride Month.
@@
Under Assumption (1), we observe $Y_{j,t}=Y_{j,t}^{0,1}$ for all $t>T^{pre}$, and we can estimate the counterfactual series $Y_{j,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of character $j$ before the beginning of the 2022 LGBT Pride Month. We can then estimate $\gamma_{t}^{L}$ by computing the differences between character $j$’s observed pick rates and the synthetic counterfactual for all $t>T^{pre}$:
@@
Under Assumption 2, the relationship $\tau_{t}^{C}=\tau_{t}-\gamma_{t}^{L}$ holds. Thus, if the estimated effects of LGBT Pride Month on players’ preferences for LGB characters are small relative to the estimated effects of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4.1 must be primarily attributed to Graves’ disclosure.
[…]
[sorting = nty]